-1/9(x-27)+1/3(x+3)=x-59
Answer by rfer(16322)

-1/9x+3+1/3x=x-59
-1/9x+3/9x-9/9x=-62
-7/9x=-62
x=-62/(-7/9)
x=-62*-9/7
x=80